OpenStudy (anonymous):

Evaluate the function f(x)=4*7^x for x = -1 and x = 2. Show your work

OpenStudy (anonymous):

You just insert x =-1 and x=2 in the function, so you get:\[f(-1)=4*7^{-1}\] and \[f(2)=4*7^2\]